In the past, Ben has been able to hit a baseball at the batting cage 20% of the time. He decides to head back to the batting cage. (a) What is the probability that his first hit will be on the eighth pitch? (b) What is the probability that his first hit will not be within the first nine pitches?
Answer #1

(A) using geometric distribution formula

where n = 8 and p = 20/100 = 0.20

(B) probability that his first hit will not be within the first nine pitches = (1-p)^n

where p = 0.20 and n = 9

this gives us

P(no hit within 9 pitches) = (1-0.20)^9 = 0.1342